OpenStudy (anonymous):

Written in scientific notation, a number is 1.025 × 10 to the fifth power In standard form, the number is _____. 1.02500 0.00001025 102,500 102,500,000

OpenStudy (karatechopper):

That 5th power defines how many steps to the right you take if it is positive, thus move your decimal 5 places to the right to find your answer.

OpenStudy (anonymous):

That 5th power defines how many steps to the right you take if it is positive, thus move your decimal 5 places to the right to find your answer. -karatechopper 1.025(10^5) =1.025*10^5 =1.025*(10*10*10*10*10) =102500 (Your answer is C!)

OpenStudy (anonymous):

@camshaw

OpenStudy (anonymous):

Thank you so much guys. got some more questions coming up keep a lookout

OpenStudy (anonymous):

Okay, your welcome!